OpenStudy (anonymous):

(-31)+x=-15 little help ?

Mehek (mehek14):

-15 + 31 = ?

Mehek (mehek14):

you can just subtract 31 - 15

OpenStudy (dinamix):

16

Mehek (mehek14):

@dinamix no giving direct answers

OpenStudy (anonymous):

x=?

Mehek (mehek14):

you have to subtract 31 - 15 to get x

Mehek (mehek14):

or add -15 + 31 either way, you get the same answer

OpenStudy (anonymous):

thanks are you sure this is a make up test i could bring home and dont really want to fail.

Mehek (mehek14):

of course im sure I already took Algebra 1

OpenStudy (anonymous):

thanks

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (anonymous):

can you help with a few more

Mehek (mehek14):

sure

OpenStudy (anonymous):

x/2=25

Mehek (mehek14):

multiply 25 * 2 = 50 x = 50

OpenStudy (anonymous):

ah i was close divided by a negetive

OpenStudy (anonymous):

(26) + x = -3

Mehek (mehek14):

-3 - 26 = -29 = x

OpenStudy (anonymous):

now what i add like terms right?

Mehek (mehek14):

subtract -3 - 26 = -29 x = -29

OpenStudy (anonymous):

thanks man x - 13 = -12 i get rid of the -13 first by adding

Mehek (mehek14):

-12 + 13

OpenStudy (anonymous):

x=1

Mehek (mehek14):

correct

OpenStudy (anonymous):

36 + x = 27

OpenStudy (anonymous):

36 - 27?

Mehek (mehek14):

27 - 36

OpenStudy (anonymous):

you reverse it so

OpenStudy (anonymous):

x = - 9

Mehek (mehek14):

correct

OpenStudy (anonymous):

x/-6 = =6

Mehek (mehek14):

6*-6

OpenStudy (anonymous):

no soultion ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1 or 0

Mehek (mehek14):

-36

Mehek (mehek14):

you aren't adding it's multiplication

OpenStudy (anonymous):

ah thx so much thats it medal and fan

Mehek (mehek14):

yw ^_^

OpenStudy (aj_harleyy):

d